Circular No. 7485

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

MXB 1659–29 R. Wijnands, Massachusetts Institute of Technology; T. Strohmayer,

Goddard Space Flight Center; and L. M. Franco, University of Chicago, report: “Using archival data from the PCA instrument on the Rossi Xray Timing Explorer, we have detected x-ray brightness oscillations with a frequency of ∼ 567 Hz in five type-I x-ray bursts observed from the x-ray transient and eclipsing binary MXB 1659–29. This source is still active more than a year after its reappearance in Apr. 1999 (IAUC 7138, 7139). The oscillations have fractional rms amplitudes of 10–20 percent and are most likely related to the spin frequency of the neutron star in MXB 1659–29. The oscillations are preferentially detected between orbital phase 0.05 and 0.75 (using the orbital ephemeris provided by Wachter et al. 2000, Ap.J. 534, 367) and only marginally detectable during intervals of x-ray dipping activity. This strongly suggests that the mechanism behind the x-ray dips inhibits the detectability of the oscillations. More observations are needed of MXB 1659–29 outside the x-ray dips and x-ray eclipses (i.e., between orbital phase 0.05 and 0.75) in order to study these oscillations before the source returns to quiescence.”

COMETS C/1996 E2, C/1997 K5, C/1997 K6, C/1998 F2 (SOHO) Further to IAUC 7480, D. Hammer and D. Biesecker have provided

measurements of four more Kreutz sungrazing comets detected by SOHO instruments; discovery positions and citations to the MPECs containing the reduced measurements and orbits by B. G. Marsden are given below. The 1997 objects (which both showed obvious tails) were found by M. Oates in archival images, and the other two comets were found by M. Meyer. All four comets were seen with the C2 coronagraph, but only C/1996 E2 (which showed cometary appearance and a tail) was also seen with C3.

Comet  1996 UT  α_{2000}  δ_{2000}  MPEC
C/1996 E2 Mar. 10.281  23^{h}44.^{m}2  − 4^{o}48′  2000-Q18
Comet  1997 UT  α_{2000}  δ_{2000}  MPEC
C/1997 K5 May 20.264  3^{h}46.^{m}9  +18^{o}14′  2000-Q18
C/1997 K6  31.188  4 31.0  +20 34  2000-Q18
Comet  1998 UT  α_{2000}  δ_{2000}  MPEC
C/1998 F2 Mar. 27.879  0^{h}33.^{m}5  + 2^{o}06′  2000-Q35
2000 September 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green